### v2

# Quantum algorithms in distributed quantum computing

Abstract
Distributed quantum computing (DQC), provides a way to scale quantum computers using multiple quantum processing units (QPU) which are connected through quantum communication links. In this paper, we have built a distributed quantum computing simulator and used the simulator to investigate quantum algorithms such as the quantum Fourier transform, quantum phase estimation, quantum amplitude estimation, and the generation of probability distribution in DQC. The simulator can be used to generate and execute distributed quantum circuits with ease, obtain and benchmark DQC parameters such as the fidelity of the algorithm and the number of entanglement generation steps, and use dynamic circuits in a distributed setting to improve results. We show the applicability of dynamic quantum circuits in DQC, where mid-circuit measurements, local operations, and classical communication are used in place of noisy inter-processor (nonlocal) quantum gates.
1 Introduction
To make quantum computers advantageous in industry, the scalability of the hardware needs to be significantly increased [Buhrman2003, Beals2013, Ferrari2021, Parekh2021]. Increasing the number of qubits increases the crosstalk, thereby increasing the gate and measurement error rates [PhysRevA.104.052404]. Distributed quantum computing (DQC) [Davis2023, Gyongyosi2021] offers an interesting approach. Instead of scaling up, we use several quantum processing units (QPUs) with fewer qubits connected through quantum communication links to increase computational capabilities [Lim2005, Niu2023]. The nodes of the DQC (Fig. 1) consist of the local qubits necessary for computation, and communication qubits [PhysRevLett.118.250502], which are used to entangle the nodes and apply quantum gates between them [Eisert2000, Monroe2009, Krutyanskiy2019].
Figure 1: Schema for the DQC, a) Showing multiple quantum processing units. b) A two-node quantum computer showing entangled (dark blue) communication qubits and local (blue) processing qubits
The communication qubits can be entangled using a probabilistic entanglement generation procedure that relies on successful detector click patterns [Nature2013, PhysRevLett.124.110501, PhysRevLett.120.200501]. The generation process is repeated until successful click patterns are obtained, indicating that the qubits have been entangled. Entanglement generation between remote nodes has been experimentally realized in superconducting qubits [PhysRevLett.120.200501], ions [PhysRevLett.124.110501], and NV centers [Nature2013] respectively. We will briefly study the success probability of this procedure.
Suppose the total efficiency (including qubit-node coupling) of the channel is $p$. For $k$ repetitions (trials), the probability of obtaining at least one entangled pair with 1 communication qubit/node is given by $P=(1-{(1-p)}^{k})$. To increase the probability of success, we can have multiple communication ($m$) qubits per node, where the entanglement generation procedure is run for each pair in parallel. The probability that we have at least one entangled pair is given by, $P_{s}=1-(1-P)^{m}$. The entangled pair is used to apply a remote quantum gate between local qubits in the different nodes utilising one ebit/CNOT gate by applying the quantum circuit shown in Fig. 2 [Eisert2000].
[…]
Figure 2: The quantum circuit for the non-local teleportation-based CNOT gate between $q_{0}$ and $q_{3}$ consists of the following steps: a) The initial quantum state of $q_{1}$ and $q_{2}$ is prepared in the Bell state $\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle$), b) two local CNOT gates, followed by measurement on the qubits $q_{1}$ and $q_{2}$, classical communication and conditional single-qubit quantum gates
The non-local CNOT gate has been experimentally realized in superconducting qubits [Chou2018] and ion qubits [Daiss2021]. A non-local quantum circuit could be obtained using multiple non-local quantum gates, where the gates are applied sequentially through entangling, measuring, and re-entangling the same qubit with only one extra qubit/node.
In this paper, we describe a distributed quantum computing simulator (DQCS) that uses non-local quantum gates between the nodes, creates distributed quantum circuits, and evaluates their performance in the presence of noise [Eisert2000]. We use the DQCS to study distributed quantum algorithms [Zhou2023, Cirac1999, Yimsiriwattana2004], demonstrate the applicability of dynamic quantum circuits in DQC that reduce resource overhead, and achieve high fidelity. We illustrate this with an example of distributed quantum phase estimation (DQPE). We study the probability distribution loading in DQC, scaling with the number of nodes. We use the DQCS to run the distributed quantum amplitude estimation (DQAE) algorithm on multiple nodes. The DQCS contains modules that provide users various ways to study, run, and evaluate the performance of quantum algorithms in DQC [Cirac1999]. The DQCS could be employed for simulating and optimizing both the quantum network layer and the quantum circuit layer for distributed quantum algorithms.
Several distributed quantum computing [Ferrari2021, Rhea2021, Hanner2021, Cuomo2020] and quantum internet simulators [Dahlberg_2019, Dahlberg_2022, QUISP] that have been developed that consider distributed quantum gates, compilers, and classical networks respectively. The goal of the DQCS is not only to consider the realistic setting for DQC, including node efficiency, gate imperfections, and dynamic quantum circuits to improve fidelity, which are crucial for experiments, but also to investigate and benchmark the performance of quantum algorithms in DQC.
2 Noise model for non-local quantum gates
The noise model for non-local quantum gates can be studied using the depolarization channel $E$ that acts on the quantum state of $n$ qubits $\rho$, $\rho\rightarrow E(\rho)$ as,
$E(\rho)=(1-\lambda)\rho+\lambda Tr(\rho)\otimes\frac{I_{2^{n}}}{2^{n}},$ (1)
where $\lambda$ is the depolarization rate, $Tr(\rho)$ is the partial trace of $n$ qubits, and $I_{2^{n}}$ is the identity $2^{n}\times 2^{n}$ matrix, respectively. The single-qubit gate error channel $E$ can be obtained using the depolarization channel for $n=1$. Here, $\rho$ represents the quantum state before applying the single-qubit quantum gate ($U$) on the $i$th qubit. The channel gives $U_{i}\rho U_{i}^{\dagger}\rightarrow\rho^{\prime}$,
$\rho^{\prime}=(1-\epsilon_{d})U_{i}\rho U_{i}^{\dagger}+\epsilon_{d}Tr_{i}(\rho)\otimes\frac{I_{2}}{2},$ (2)
where $\epsilon_{d}$ is considered as the depolarization single-qubit gate error rate (respectively) and $Tr_{i}$ is the partial trace of the $i^{th}$ qubit. Let us consider a two-qubit quantum gate $U_{i,j}$ (between the qubits $i$ and $j$). The two-qubit quantum gate error channel could be obtained for $n=2$ that applies the transformation $U_{ij}\rho U_{ij}^{\dagger}\rightarrow\rho^{\prime\prime}$.
$\rho^{\prime\prime}=(1-\epsilon_{g})U_{ij}\rho U_{ij}^{\dagger}+\epsilon_{g}Tr_{i,j}(\rho)\otimes\frac{I_{4}}{4}$ (3)
where $\epsilon_{g}$ is the two-qubit quantum gate error rate and $Tr_{i,j}$ is the partial trace of the qubits $(i,j)$. Note that in the case of a local CNOT gate, the fidelity from the error model is $\epsilon_{g}/2$. The fidelity of the quantum circuit after one successful implementation of the non-local CNOT gate (up to the first order) is $\epsilon\approx(\epsilon_{d}+\epsilon_{g})$. If we consider $\epsilon_{m}$ to denote the measurement reset error, we may approximate that the fidelity for one distributed quantum gate is $(1-(\epsilon_{d}+\epsilon_{g}+2\epsilon_{m}))$. For $N$ distributed quantum gates,
$F\approx{(1-(\epsilon_{d}+\epsilon_{g}+2\epsilon_{m}))}^{N}$ (4)
3 Distributed quantum computing simulator
We built a distributed quantum computing simulator (DQCS) [GitHub] that automates the creation of distributed quantum circuits for distributed quantum algorithms. It consists of the following modules:
  • Non-local quantum gates: This module facilitates the application of non-local quantum gates between qubits residing in different nodes. It operates the quantum circuit in Fig. 2, which enables the quantum circuit to span across nodes.
  • Distributed circuits: The distributed circuits module enables users to effortlessly create distributed quantum circuits using a node dictionary and a quantum circuit. It automatically handles communication registers, so users don’t need to explicitly specify them.
  • Create distributed circuits: To create the distributed quantum circuits for advanced quantum algorithms from existing quantum circuits, this module offers flexibility where the user can provide input to the communication register and obtain the distributed quantum circuit.
  • Noise model: We can experiment with the distributed quantum circuit in the presence of noise, the library offers an optional noise model module. This component incorporates realistic noise modeling, including single-qubit and two-qubit errors. The noise parameters can be changed corresponding to different experimental systems and the fidelity of the quantum algorithm can be obtained. Note that in place of the entanglement generation, the Bell pair is created using a CNOT gate in the simulator. Hence, there is a small difference in the noise model where the depolarization of the Bell pair is approximated with the gate error $\epsilon_{g}/2$.
The steps to run the simulator are as follows:
  1. 1.
    Create a quantum circuit and instantiate the distributed circuits with your quantum circuit and the nodes to obtain the distributed quantum circuit.
    import DQCS
    qc = QuantumCircuit(3)
    qc.h(0)
    qc.h(1)
    qc.cx(0, 2)
    nodes = {"1": [0, 1], "2": [2]}
    gate_app, circuit = DQCS.DistributedCircuits(qc, nodes).create_circuit()
    The circuit gives the distributed quantum circuit the parameter $gate\textunderscore app=1$ upon the successful creation of the distributed quantum circuit.
  2. 2.
    Alternatively, use the $CreateDistributedCircuits$ instance with the quantum circuit, communication qubits, nodes, and the probability $p$ (coupling efficiency for the node) to create the distributed quantum circuit. Let us consider the example to illustrate the module.
    import DQCS
    q_0 = QuantumRegister(2, ’a’)
    q_1 = QuantumRegister(2, ’b’)
    qc = QuantumCircuit(q_0, q_1)
    nodes = {"1": [q_0[0], q_0[1]], "2": [q_1[0]], "3": [q_1[1]]}
    qc.h(q_0[0])
    qc.cx(q_0[1], q_1[1])
    #Define an empty circuit with communication registers
    comm = QuantumRegister(4, ’c’)
    c = ClassicalRegister(4, ’cl’)
    circ = QuantumCircuit(q_0, q_1, comm, c)
    gate_app, circuit = DQCS.CreateDistributedCircuits(qc, c, circ, comm, nodes, 1).create_circuit()
    The probability $p=1$ in the above example is the qubit-node coupling efficiency. This module is more suitable for running Qiskit’s in-built quantum algorithms.
  3. 3.
    Instantiate the noise model with your distributed quantum circuit, specifying parameters such as the number of shots, repetitions, and error probabilities (single, CNOT quantum gates, measurement, reset errors). We show an example of including a noisy $CNOT$ gate in the simulation.
    noise_model = noise.NoiseModel()
    noise_model.add_all_qubit_quantum_error
    (error_gate2, [’cx’])
    basis_gates = noise_model.basis_gates
    hist = execute(qc, simulator,basis_gates,noise_model, shots).result()
4 Distributed quantum algorithms
Quantum algorithms are utilized to achieve speedup for various applications including machine learning, optimization [Farhi2014], credit risk analysis [Egger2021], and option pricing using Monte Carlo integration [Nik2020, Carrera2021, Herbert2022]. We consider the simulation of two distributed quantum algorithms, DQPE and DQAE, in DQC which are the underlying quantum algorithms for more advanced quantum algorithms [Egger2021, Nik2020, Carrera2021, Herbert2022].
4.1 Dynamic quantum circuits
Dynamic quantum circuits involve the use of mid-circuit measurements, qubit reset, reuse, and single-qubit conditional quantum gates [PhysRevLett.127.100501]. Here, we explore the use of dynamic quantum circuits instead of distributed quantum gates in the DQC [dynamic]. One of the examples of dynamic quantum circuits is when we have a quantum Fourier transform followed by measurement [dynamic].
Let’s consider a quantum circuit for the QFT on three qubits in Fig. 3, followed by measurement of the qubits.
[…]
Figure 3: Quantum Fourier transform on three qubits, where $P(\theta)$ refers to the phase gate
The corresponding dynamic quantum circuit is shown in Fig. 4.
[…]
Figure 4: Dynamic quantum Fourier transform on three qubits
We can now illustrate the application of distributed dynamic QFT to the quantum phase estimation.
4.2 Distributed dynamic quantum phase estimation
The problem of quantum phase estimation is to determine the eigenvalue of the unitary gate that acts as $U|\psi\rangle=e^{2i\pi y}|\psi\rangle$, where $|\psi\rangle$ is the eigenstate of $U$, and $y$ is a fraction corresponding to the eigenvalue. The quantum circuit used to implement QPE is shown in Fig. 5.
[…]
Figure 5: Distributed dynamic quantum phase estimation, where the qubits are in different nodes. The controlled unitary gates are applied using DQCS, followed by dynamic QFT
We can derive the fraction $y$ from the outcome of the measurement. For instance, if we obtain the measurement outcome $y_{0}y_{1}y_{2}$, where each $y_{i}\in{0,1}$, then the fraction $y$ is calculated as $(y_{0}y_{1}y_{2})_{10}/2^{3}$. To illustrate, let’s examine the implementation of Quantum Phase Estimation (QPE) for the unitary gate $U=R_{X}(\pi)\otimes R_{Y}(\pi)$ acting on two qubits. One of the eigenvalues for this operator is $-1$ ($y=1/2$), corresponding to a measurement output of $100$. We prepare the eigenstate $|\psi\rangle$, create the distributed QPE (DQPE) for the following qubit distributions: 1: [0, 1, 2], 2: [3, 4], and apply the dynamic QFT to the distributed quantum circuit. The probability of the outcome is shown in Fig. 6, where we consider the locally controlled unitary gates and quantum Fourier transform for the QPE (dark blue) (Fig. 3). For the DQPE (blue), we consider distributed two-qubit quantum gates and dynamic QFT without two-qubit quantum gates for the quantum Fourier transform (Fig. 4). From the noise model, the dynamic quantum Fourier transform doesn’t suffer from two-qubit quantum gate errors, only measurement and depolarization errors, giving better results in Fig. 6. For quantum phase estimation using $n$ control qubits, we need $N$ two-qubit quantum gates to realize the controlled-unitary operations and the quantum Fourier transform requires $n(n-1)/2$ two-qubit quantum gates. Assuming the same error for the two-qubit (controlled-unitary and controlled phase) gates, the fidelity of the algorithm is approximated $F\approx(1-\epsilon_{g})^{N+\frac{n}{2}(n-1)}$. For the DQPE, where we have $N$ distributed quantum gates (each applied non-locally using 2 CNOT quantum gates), the fidelity is given by $F\approx(1-\epsilon_{g})^{2N}$.
Figure 6: The probability of the outcome for the distributed QPE (DQPE) (blue) vs. QPE (red) in the presence of noise $\epsilon_{d}=\epsilon_{g}=\epsilon$, $\epsilon_{m}=0$
4.3 Distributed quantum amplitude estimation
Quantum amplitude estimation (QAE) is used to calculate the unknown amplitude of a quantum state. For a given unitary operator $A$,
$A|0\rangle_{n+1}=\sqrt{1-a}|\psi_{0}\rangle_{n}|0\rangle+\sqrt{a}|\psi_{1}\rangle_{n}|1\rangle,$ (5)
where $a$ is the unknown amplitude to be estimated, and $|\psi_{0}\rangle_{n}$ and $|\psi_{1}\rangle_{n}$ are (normalized) good and bad states, respectively, that correspond to the probability of measurement of 00$\rangle$0 ⟩ and $|1\rangle$ in the last qubit, respectively.
QAE has applications such as calculating the expectation value of the discrete linear function, quantum Monte Carlo integration [Nik2020, Herbert2022], quantum risk analysis, and portfolio optimization. We use the DQCS to simulate distributed QAE (DQAE) using multiple nodes. There are multiple approaches to performing QAE. Here, we study DQAE using maximum likelihood estimation [Suzuki2020], which achieves quadratic speedup over classical Monte Carlo algorithms in DQC without using the QFT. We show that the estimation error for the DQAE [Suzuki2020], simulated using the DQCS, achieves the Cramer-Rao bound, which provides a way to scale QAE using multiple QPUs.
We prepare the $A$ operator, using the probability distribution loading $P(x)$ [Nik2020], that loads a discrete probability distribution to a quantum state. The bounded function $f(x)$ is constructed from the samples of $P(x)$ and realized using multiple controlled-$R_{y}$ gates with controlled-$R_{y}(f(i))$ on the gate on the $ith$ qubit [Nik2020]. The probability of the measurement of $|1\rangle$ on the last qubit gives the expectation value $\mathrm{E}[f(x)]$.
$\mathrm{E}[f(x)]=\sum_{i=0}^{2^{n}-1}p(x_{i})f(x_{i}),$ (6)
Suppose P(x) is a uniform distribution, the probability of measuring $|1\rangle$ on the last qubit is given by $\frac{1}{2^{n}}\sum_{i=0}^{2^{n}-1}f(x_{i})$ for $x_{i}=i/2^{n}$ [Carrera2021]. For QAE, we construct the $Q$ operator, such that $Q=-AS_{0}A^{-1}S_{X}$. The operator $S_{X}$ leaves the good states unchanged and includes a phase $(-1)$ on the bad states. $S_{X}|\psi_{0}\rangle|0\rangle=|\psi_{0}\rangle|0\rangle$, $S_{X}|\psi_{1}\rangle|1\rangle=-|\psi_{1}\rangle|1\rangle$. The operator $S_{0}$ attaches a phase (-1) to $|0\rangle_{n}$ and leaves other basis states unchanged, the repeated application of $Q$, $M$ times, on the state $A|0\rangle$ gives
$Q^{m}|\psi\rangle=\cos(2M+1)\theta|\psi_{0}\rangle|0\rangle+\sin(2M+1)\theta|\psi_{1}\rangle|1\rangle,$ (7)
QAE requires multiple controlled unitary gates for the implementation of the reflection operators. We could write multiple CNOT gates (with $n$ control qubits and a single-qubit unitary) with single-qubit gates and $\leq(20n-38)$ CNOT gates [Rv2023, Eisert2000]. We consider the QAE without phase estimation that uses the maximum likelihood estimate [Suzuki2020] for the integration of $\int_{0}^{c}\sin^{2}(x)$ in DQC. We compare the estimated error values against the Cramer-Rao bound [Suzuki2020], where $P(x)$ is a uniform distribution obtained by applying Hadamard gates to all of the qubits. The function $f(x)$ is prepared with $a=c/2^{n-1},b=c/2^{n}$ using 1 $R_{y}$ gate and $n$ controlled-Ry gates, $Ry(c/2^{n-i})$ [Suzuki2020]. It can be seen in Fig. 7 (($c=\frac{\pi}{3}$)) that DQAE can be simulated and the estimated value achieves the Cramer-Rao bound.
Figure 7: The estimated value in a quantum computer (1 node), distributed quantum computer (2 nodes), and Cramer-Rao bound [Suzuki2020]
In our simulation, we used Qiskit’s transpile function to apply a distributed multiple controlled gate and distributed controlled unitary gates using non-local CNOT gates (Fig. 2). The DQCS can be used to obtain the resource estimation for the DQAE for various node parameters in Table 1.
$\text{n}_{\text{nodes}}$ $\text{n}_{\text{Grover}}$ $k$ Estimation  error
(2,3) 00 [(6,7,9),(6,7,11)] (0.00881913, 0.01116442)
(2,3) $1$ [(28,30,48),(30,34,54)] (0.0093686, 0.02658211)
(2,3) $2$ [(50,59,91),(54,64,101)] (0.00576238, 0.00911406)
(2,3) $4$ [(94,119,157),(102,113,193)] (0.00498004, 0.0015308)
(2,3) $8$ [(182,220,323),(198,238,338)] (0.00251507, 0.00070526)
(2,3) $16$ [(358,441,687),(390,449,758)] (0.0011292, 0.00141048)
(2,3) $32$ [(710,900,1366),(774,908,1371)] (0.00050303, 0.00017524)
Table 1: Resource estimation for the distributed quantum amplitude estimation (DQAE) obtained using the DQCS where the number of nodes $(n)$, number of Grover operators $(m)$, number of entanglement generation steps $k$ for the qubit-node coupling parameter $p=(1,0.8,0.5)$, and estimation error for $m=1$
are considered
We could also consider QAE using normal distribution instead of uniform distribution. It is crucial to understand the quantum state preparation of probability distribution in DQC. The normal distribution
$P(X=x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{\frac{-(x-\mu)^{2}}{\sigma^{2}}},$ (8)
loaded to quantum state $|p\rangle=\sum_{i=0}^{2^{d-1}}\sqrt{p_{i}}|x_{i}\rangle$, where $p_{i}$ is sampled from $P(x)$ with $2^{d}$ points [Nik2020, Carrera2021]. We use the DQCS to simulate the quantum state preparation of a probability distribution. We consider a normal distribution with $8$ qubits with the mean $\mu=1$ and $\sigma=2$ in the DQCS, with $n=1,2,4,8$ nodes, to obtain the Heilinger fidelity in Fig. 8.
Figure 8: The Heilinger fidelity of a normal distribution with 8 qubits with different numbers of nodes, where $\epsilon=(0.002(\text{blue}),0.005(\text{orange}),0.009(\text{green}))$
5 Conclusion
We outlined a distributed quantum circuit simulator (DQCS), which uses a number of interconnected quantum processing units (QPUs) consisting of local and communication qubits. We studied foundational quantum algorithms such as the quantum Fourier transform, quantum phase estimation, and amplitude estimation using the DQCS and showed the application of dynamic quantum circuits in DQC. Recently, there has been growing interest in the realization of quantum networks and distributed quantum computing in various platforms such as superconducting qubits, NV centers [Nature2013, NV], ions [PhysRevLett.124.110501], and neutral atoms [neutral]. The DQCS provides a platform for users to input the experiment parameters, simulate distributed quantum circuits, and obtain the estimates for quantum algorithms including quantum phase estimation, quantum amplitude estimation, quantum state preparation of the probability distribution, etc. In the future, we aim to integrate quantum internet simulators with the DQCS to study distributed quantum algorithms in the quantum internet. We look forward to including quantum error correction and fault tolerance in the DQCS to obtain favourable polynomial scaling for the fidelity of quantum algorithms [Muralidharan2016]. We could also study distributed quantum machine learning algorithms for classification and regression [Tang2023].
Acknowledgements
I would like to thank Constantin Gonciulea, Charlee Stefanski, Vanio Markov and David Novak for helpful discussions, comments on the DQCS, and suggestions to improve the quality of the manuscript.
[…]